The perimeter of a rectangular garden is 40 feet. If the length is 6 feet more than the width, what are the length and width of the garden?
So far I have. P = 2(l+w); P = 2(6+w)+w); 40 = 2(6+w)+w)
I don't know where to go from here.

Answer by mathman409(67)

23=L 17=W
P=40
L=W+6
w+l=40
now convert l into w
w+(w+6)=40
2w+6=40
2w=40-6
2w=34
w=17
l=23 because 17(or w)+6=23 which is l